OpenStudy (marigirl):

Is it possible to model a Cubic so that it has a maximum of (30,12). See my drawing below

OpenStudy (marigirl):



OpenStudy (solomonzelman):

Cubic function doesn't have an absolute max or absolute minimum if that is what you are referring to.

OpenStudy (solomonzelman):

We can model a local maximum at (30,12) if you meant that.

OpenStudy (marigirl):

use the point (30,12) and find 𝑦=1/6750 𝑥(𝑥−60)(𝑥−120),

OpenStudy (marigirl):

but still it is not the maximum......... prob cuz it doesnt have it (like you said)

OpenStudy (solomonzelman):

here: desmos.com use that calculator to graph your equation, and you will see that (30,12) is not the local maximum (in fact not even a point on the function).

OpenStudy (marigirl):

thanks :)

OpenStudy (solomonzelman):

You can do this: \(y=\left(x-30\right)^3+12\) (Shift a parent function x³ by 30 units to the right, and by 12 units up)

OpenStudy (solomonzelman):

Doesn't that function seem to have a local maximum at (30,12)?

OpenStudy (solomonzelman):

Or even better, \(y=\left(x-30\right)^3-\left(x-30\right)^2+12\)

OpenStudy (solomonzelman):

the second part I added, fixes the chape of the local maximum.

OpenStudy (marigirl):

great. thanks heaps. I really appreciate it!

OpenStudy (solomonzelman):

Sure, yw